November 18, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, November 18, 2015
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at .

Calibration Tasks


  1. Announcements
  2. Calibration Data Taking
  3. Calibration Tasks
  4. Subdetector Reports
  5. Future Simulations
  6. AOB


Attending: Sean (NU); Simon, Eugene, Elton, Mark I., Justin, Nathan, Alex B., Luke, Lubomir, Paul M. (JLab); Curtis, Mike S., Will M., Naomi (CMU); Mahmoud (FIU); Brad (FSU)

  1. Announcements
    • Still on track for beam on Dec. 4
    • The lustre-based disks available at the JLab farm (e.g. work and volatile disks) were acting up this morning. The latest word is that fixes are being implemented and
  2. Calibration Data Taking
    • Adesh gave Sean some updated numbers for the FCAL - at least 10 times as much data is needed.
  3. Calibration Tasks
    • Sean briefly discussed the meeting last Friday for the Calibration Challenge. Most of the discussion was about fitting the different pieces together that people are working on. Additions to sim-recon have been coming in - all parties expect to have their code committed by the end of the week.
    • We'll meet again on Friday, pending sim-recon and farm status
  4. Subdetector Reports
    • FCAL - Adesh has generated updated pedestal and gain calibrations and will upload them to the CCDB. He will work with Manuel to do the hardware gain adjustments. Next on his list is performing the timing calibrations and getting these codes into sim-recon.
      • Later in the meeting Eugene asked a couple of questions:
      • Q: How close are we to design resolution? A: current pi0 width is 8.7 MeV, compared to the 7 MeV expected.
      • Q: How well calibrated is the periphery compared to the center of the detector? A: Hasn't been studied, will look at that.
    • BCAL - Not much news - Mike still needs to finish checking his reorganization of the BCAL timing. He expects to have that done today.
    • CDC - Mike has obtained improved time-to-distance relations by fitting data, which give resolutions close to design values. When applying these, cosmic data looks good, but beam data still has some features that need to be understood to finalize this project.
    • FDC - Lubomir reported that he has obtained improved resolutions (~150 um) with several improvements to the reconstruction, including improved fitting of the strip clusters. Simon continues to work to incorporate these into the standard tracking code.
    • TOF - Mark reported that Sasha from FSU is trying to take cosmic data to do gain calibrations, but is finding few good events. An issue with the TDC timing offset that made the timing of every 6 events was found and fixed - Mike uploaded new timing calibrations as of yesterday.
    • Start Counter - Mahmoud successfully tested his new timewalk code, and will add it to sim-recon today. He finds a timing resolution of 270 ps. He will then work on converting the propogation time and attenuation length calibrations.
    • TAGH - Nathan is working on getting his timewalk calibration code into sim-recon. He's currently cleaning up his scripts and hopes to have it finished today.
    • TAGM - Alex is working on updating the PS and TAGM timewalk codes and getting them into sim-recon. He's also planning to add his PS energy calibrations. He hopes to have these in by the end of the week.
    • RF - Paul has updated the RF calibration routines.
  5. Future Simulations
    • Mark submitted 1000 test jobs for the 2016 simulation conditions. Sean looked at these and pronounced them "good".
    • Mark has added in some monitoring plugins and is testing these.
    • Sean wrote some code to save BCAL and FCAL energy sums with per-channel thresholds similar to what the trigger uses, so users can study the effects of cuts on these quantities. He will push the code to the master branch today. Paul said he was okay with this idea.
    • Once these changes are validated, we can begin production.